Calculate A×B Where A is 4m and B is 3m for two vectors and the angle between them is 450



1
Expert's answer
2021-02-23T10:03:21-0500

The vector product of two vectors can be found as follows:


"\vec{A}\times\vec{B}=|\vec{A}|\cdot|\vec{B}|sin\theta,""\vec{A}\times\vec{B}=4\ m\cdot3\ m\cdot sin45^{\circ}=8.48\ m^2."